OpenStudy (anonymous):

A guitar string with mass density μ = 2.1 × 10-4 kg/m is L = 1.09 m long on the guitar. The string is tuned by adjusting the tension to T = 108.7 N. How do I find the speed of the waves on the string?

OpenStudy (anonymous):

Good. We have the tension. The velocity is v = sqrt(T/mu), where T is tension (force) and mu is mass per unit length of the string. v=sqrt(108.7N/2.1x10-4kg/m)=720 m/s